OpenStudy (anonymous):

WILL FAN AND MEDAL In the formula rt=d r is rate t is time d is distance what is the rate for a runner that completes 7.5 miles in 1.5 hours a 11.25 b9.0 c6.0 d5.0

OpenStudy (idealist10):

If rt=d and you want to find the rate, then r=d/t, so r=7.5/1.5=5. Therefore, the answer is D.